OpenStudy (anonymous):

how would i answer this problem? Angela currently has a account balance of $4,298.55. she opened the account 15 years ago with a deposit of $1,987.22. If the interest compounds monthly, what is the interest rate on the account? I know it looks like this when plugged into the formula: 4,298.55= 1,987.22(1+x/12)^12*15 how do i solve this though?

jimthompson5910 (jim_thompson5910):

You have the right values and you plugged them in the right way. So far so good

jimthompson5910 (jim_thompson5910):

Let's now solve for x 4298.55= 1987.22(1+x/12)^(12*15) 4298.55= 1987.22(1+x/12)^(180) 4298.55/1987.22 = (1+x/12)^(180) 2.16309719105081 = (1+x/12)^(180) (2.16309719105081)^(1/180) = 1+x/12 1.00429553882183 = 1+x/12 1.00429553882183-1 = x/12 0.00429553882183 = x/12 0.00429553882183*12 = x 0.05154646586196 = x x = 0.05154646586196 So the interest rate is roughly 5.15%

OpenStudy (anonymous):

so i dont have to use ln?

jimthompson5910 (jim_thompson5910):

no because there isn't a variable in the exponent

OpenStudy (anonymous):

oh ok thanks

jimthompson5910 (jim_thompson5910):

yw